# BeyondMimic: From Motion Tracking to Versatile Humanoid Control via Guided Diffusion

### v2

@@ Abstract
Learning skills from human motions offers a promising path toward generalizable policies for whole-body humanoid control, yet two key cornerstones are missing: (1) a high-quality motion tracking framework that faithfully transforms large-scale kinematic references into robust and extremely dynamic motions on real hardware, and (2) a distillation approach that can effectively learn these motion primitives and compose them to solve downstream tasks. We address these gaps with BeyondMimic , a real-world framework to learn from human motions for versatile and naturalistic humanoid control via guided diffusion. Our framework provides a motion tracking pipeline capable of challenging skills such as jumping spins, sprinting, and cartwheels with state-of-the-art motion quality. Moving beyond mimicking existing motions and synthesize novel ones, we further introduce a unified diffusion policy that enables zero-shot task-specific control at test time using simple cost functions. Deployed on hardware, BeyondMimic performs diverse tasks at test time, including waypoint navigation, joystick teleoperation, and obstacle avoidance, bridging sim-to-real motion tracking and flexible synthesis of human motion primitives for whole-body control. https://beyondmimic.github.io/.
@@ I Introduction
Recent advances in physics-based character animation [1, 2, 3, 4] have already shown incredible progress to synthesize human motions into dynamic behaviors for whole-body control to directly solve downstream tasks. However, these successes remain limited to simulation, where agents enjoy idealized dynamics, unlimited actuation, and perfect observations, unlike real-world humanoids that face unmodeled dynamics, hardware limits, and imperfect state estimation. Achieving similar capabilities on hardware requires two missing capabilities: (1) a scalable, high-quality motion tracking framework that transforms kinematic references into robust, highly-dynamic motions while avoiding over-randomization, jitter, and degraded motion quality seen in prior work, and (2) an effective sim-to-real recipe to distill learned motion primitives into a single policy capable of composing skills at test time for flexible, goal-driven control without retraining.
@@
We introduce BeyondMimic, a real-world framework to address the above two challenges altogether. BeyondMimic first introduces a motion tracking pipeline that is able to perform highly dynamic motions, such as jumping spins, sprinting, and cartwheels, on real humanoid hardware with state-of-the-art motion quality. Beyond simply mimicking these human motions on humanoids, our framework further presents a guided diffusion policy that synthesizes these motions or skills into a single unified policy that supports zero-shot, task-specific control at test time using only simple cost functions, solving downstream tasks flexibly without additional retraining.
@@
  • Guided Diffusion for Humanoids: Application of loss-guided diffusion to real-world humanoid whole-body control, enabling diverse downstream tasks via simple loss functions at test time.
@@ II-A Motion Tracking
Earlier motion-tracking approaches typically rely on a small set of similar reference motions to train single-task policies. These policies are co-trained with task-specific rewards alongside the DeepMimic objective, producing controllers capable of performing one task in one style. Examples include quadrupedal locomotion [10] and goalkeeping [11], and bipedal jumping and running [12]. To enable more dynamic skills, ASAP [13] proposes to use a real-to-sim pipeline that learns a delta action model from hardware experiments, improving simulation fidelity. However, this approach requires training motion-specific delta action models, which tend to overfit to the exact short motion. Similarly, KungfuBot [14] and HuB [15] achieve high-quality sim-to-real transfer with carefully designed domain randomization, but only for single motions.
To move beyond single-motion policies, recent research explores scalable motion-tracking frameworks that can learn a diverse set of motions within a single policy. In physics-based animation, PHC [16] is a leading example of such a system, inspiring subsequent attempts in robotics to build general-purpose motion trackers. Early multi-motion trackers in robotics, such as OmniH2O [17], Exbody [18], and HumanPlus [19], demonstrate the feasibility of this approach but suffer significant motion quality degradation compared to their graphics counterparts. More recently, TWIST [20] enables good quality motion tracking but primarily for static motions, and similarly, CLONE [21] and UniTracker [22] focus mainly on low-dynamic walking motions with responsive upper-body tracking. On the other hand, GMT [23] handles some dynamic motions but sacrifices global trajectory tracking in favor of relative velocity tracking and heavy lower-body regularization to improve robustness. Grandia et al. [24] present perhaps the highest-quality multi-motion tracking to date, but only on a small robot and without dynamic motions.
@@ II-B Diffusion in Robotics and Character Animation
To avoid the planning-control gap, another line of work focuses on learning a single, end-to-end policy that maps directly from state to an action distribution. This approach, Diffusion Policy [30], extends naturally to locomotion and character control. For example, Diffuseloco [31], and BiRoDiff [32] learn unified policies that encode a wide range of skills, achieving smooth gait transitions and robust real-world deployment for quadruped robots. Meanwhile, PDP [33] extend the framework to physically based characters, showing that it can follow text prompts, perform universal motion tracking, and capture the inherent multi-modal nature of human push recovery. Recently, HugWBC [34] adapted a diffusion policy to humanoid robots using online distillation. While effective, such policies lack a mechanism for flexible test-time guidance, such as loss-guided diffusion. This is because there is no straightforward way to compare a task goal defined in state space with an action sequence defined in PD targets or torques. As a result, conditioning on new tasks typically requires retraining.
@@ III Scaleable Motion Tracking
We propose a scalable pipeline for training motion-tracking policies. Using a single MDP configuration and a common set of hyperparameters, the pipeline produces high-quality sim-to-real policies given minutes-long references, demonstrating its scalability towards a diverse set of motions. We introduce the details below. Note that since our tracking framework does not have a history, we omit timestep $t$ in the formulation for clarity. We use the subscript m to denote quantities from the reference motion. Unless otherwise stated, all amounts in this section are expressed in the world frame.
@@ III-B Observations
We use an asymmetric actor–critic to boost training efficiency. In addition to the policy observation, the critic also receives per-body relative poses w.r.t. the anchor, $T^{-1}_{b_{\text{anchor}}}\,T_{b},\ \forall b\in\mathcal{B}$, enabling it to estimate body-wise tracking errors directly in Cartesian space.
@@ III-D Rewards
Task rewards are DeepMimic-style rewards. First, we compute error metrics for each target body $b\in\mathcal{B}_{\text{target}}$ based on the desired $(\hat{T}_{b},\hat{\mathcal{V}}_{b})$ and actual $(T_{b},\mathcal{V}_{b})$ poses and twists: $\mathbf{e}_{p,b}=\hat{\mathbf{p}}_{b}-\mathbf{p}_{b}$, $\mathbf{e}_{R,b}=\log(\hat{R}_{b}R_{b}^{\top})$, $\mathbf{e}_{v,b}=\hat{\mathbf{v}}_{b}-\mathbf{v}_{b}$, and $\mathbf{e}_{w,b}\approx\hat{\mathbf{w}}_{b}-\mathbf{w}_{b}$ assuming the orientaion error is small. Then, the mean squared errors are computed across all target bodies: $\bar{e}_{\chi}=\frac{1}{|\mathcal{B}_{\text{target}}|}\sum_{b\in\mathcal{B}_{\text{target}}}\|\mathbf{e}_{\chi,b}\|^{2},\,\chi\in\{p,R,v,w\}$. Each error metric is then normalized using a Gaussian-style exponential function: $r(\bar{e}_{\chi},\sigma)=\exp\left(-{\bar{e}_{\chi}}/{\sigma^{2}_{\chi}}\right)$, where $\sigma_{\chi}$ is a nominal error determined empirically.
The combined tracking reward is then defined as:
$r_{\text{tracking}}=\sum_{\chi\in\{p,R,v,w\}}r\!\bigl{(}\bar{e}_{\chi},\sigma_{\chi}\bigr{)}.$ (1)
For regularization, we include as few as three penalties critical for sim-to-real alignment. Among them, the joint limit penalty, $r_{\text{limit}}$, encourages joint positions to remain within the soft limits. The action rate penalty, $r_{\text{smooth}}$, encourages consecutive actions to be smooth, avoiding policies with excessive jitter. To penalize self-collision, we count the number of bodies where the self-contact force exceeds a predefined threshold as the total penalty, $r_{\text{contact}}$, over the end effector set $\mathcal{B}_{\text{ee}}\subseteq\mathcal{B}$.
@@
$r=r_{\text{target}}-\lambda_{l}r_{\text{limit}}-\lambda_{s}r_{\text{smooth}}-\lambda_{c}r_{\text{contact}}$ (2)
@@ III-F Adaptive Sampling
Since a failure is more likely caused by suboptimal actions taken shortly before termination, we apply a non-causal exponential decay kernel $k(\tau)=\gamma^{\tau}$ with decay rate $\gamma$ to weight the past nearby failures more heavily. The final sampling probability from bin s is then:
$p_{\text{s}}=\frac{\sum_{\tau=0}^{K-1}\alpha^{\tau}\,\bar{r}_{\text{s}+\tau}}{\sum_{j=1}^{\text{S}}\sum_{\tau=0}^{K-1}\alpha^{\tau}\,\bar{r}_{j+\tau}},$ (3)
where $\bar{r}_{\text{s}+\tau}$ is the smoothed failure rate of bin s. We further mix the probability $p_{\text{s}}$ with a uniform distribution to preserve coverage of easier bins and mitigate catastrophic forgetting, using $p_{\text{s}}^{\prime}=\lambda\frac{1}{B}+(1-\lambda)p_{\text{s}}$, where $\lambda$ is the uniform sampling ratio. Starting bins are then drawn from $\mathrm{Multinomial}(p_{1}^{\prime},\dots,p_{\text{S}}^{\prime})$, prioritizing challenging regions.
@@ IV-A Training
The training itself follows a standard denoising diffusion process. First, in a forward pass, we incrementally add Gaussian noise to ground-truth trajectories. A denoising network, $x_{0,\theta}$, then learns to reverse this corruption. It takes a noised trajectory $\bm{\tau}_{t}^{\bm{k}}$ as input and is trained to predict the original, clean trajectory $\hat{\bm{\tau}}_{t}$. The network’s parameters are optimized by minimizing the Mean Squared Error (MSE) loss against the ground truth:
$\mathcal{L}=\text{MSE}(x_{0,\theta}(\bm{\tau}_{t}^{\bm{k}},\bm{O}{t},\bm{k}),\bm{\tau}_{t})$ (4)
@@ IV-B Guidance
$\nabla_{\bm{\tau}}\log p(\bm{\tau}\mid\bm{\tau}^{*})=\nabla_{\bm{\tau}}\log p(\bm{\tau})+\nabla_{\bm{\tau}}\log p(\bm{\tau}^{*}\mid\bm{\tau})$ (6)
In our application, this guidance term is a differentiable, task-specific cost function, $G_{\bm{\tau}}^{c}(\bm{\tau})$. By establishing a relationship between the cost and the conditional likelihood, $p(\bm{\tau}^{*}\mid\bm{\tau})\propto\exp(-G_{\bm{\tau}}^{c}(\bm{\tau}))$, the guidance term simplifies to the negative gradient of the cost: $\nabla_{\bm{\tau}}G_{\bm{\tau}}^{c}(\bm{\tau})$.
@@ IV-C Downstream Tasks
$G_{\bm{\tau}}^{c}(\bm{\tau})=\frac{1}{2}\sum_{t^{\prime}=t}^{t+H}\|V_{xy,t^{\prime}}(\bm{\tau}_{t^{\prime}})-g_{v}\|^{2},$ (7)
where $\bm{\tau}$ is the predicted trajectory at timestep $t$, $V_{xy,t}$ pulls out the planar root velocity at timestep $t$ and $g_{v}\in\mathbb{R}^{2}$ is the goal root velocity from the controller.
@@
$G_{\bm{\tau}}^{\text{ts}}(\bm{\tau})=\sum_{t^{\prime}=t}^{t+H}(1-e^{-2d})\|P_{x}(\bm{s}_{t^{\prime}})-g_{p}\|^{2}+e^{-2d}\|V_{x,t^{\prime}}(\bm{\tau}_{t^{\prime}})\|^{2},$ (8)
@@
$G_{\bm{\tau}}^{c}(\bm{\tau})=\sum_{t^{\prime}=t}^{t+H}\sum_{b\in\mathcal{B}_{\text{c}}}B(\text{SDF}(\mathbf{P}_{b,t^{\prime}}(\bm{\tau}))-r_{i},\delta),$ (9)
@@ V-C Real-world Performance
TABLE II: Iterations to Convergence on Adaptive Sampling Ablation
@@ VI-A Data
Since the diffusion Real robotic systems exhibit action delay, the time lag between recieving an observation and the physical execution of the action. To handle this, we include action delay domain randomization during training for these se. This is necessary because the diffusion model operates under delayed actuation at test time
A key difference from the motion trackers used in Section III is the inclusion of action delay domain randomization. This is necessary because the diffusion model operates under delayed actuation at test time.
@@ VI-D State Representation Ablation
    • Local States: Cartesian positions ($\mathbb{R}^{3J}$) and linear velocities ($\mathbb{R}^{3J}$) of all joints, expressed in the character frame at each timestep.
@@
    • Local States: Joint angles ($\mathbb{R}^{J}$) and joint velocities ($\mathbb{R}^{J}$).
@@
Body-Pos State 100 80
Joint-Rot State 72 0
Table III depicts the performens of each representation for each of the two evaluated tasks. We find that the Body-Pos state representation significantly outperforms Joint-Rot across both tasks. While Joint-Rot is theoretically more Markovian, small errors in joint position estimates accumulate through the kinematic chain, making this representation more susceptible to compounding diffusion prediction errors compared to directly predicting Cartesian body poses. Furthermore, adding guidance on the Joint-Rot state joystick control fails quickly and abruptly, even after careful tuning, likely because the base controller lacks robustness, and guidance pushes it further out of distribution.

### v3

@@ Abstract
Learning skills from human motions offers a promising path toward generalizable policies for versatile humanoid whole-body control, yet two key cornerstones are missing: (1) a high-quality motion tracking framework that faithfully transforms large-scale kinematic references into robust and extremely dynamic motions on real hardware, and (2) a distillation approach that can effectively learn these motion primitives and compose them to solve downstream tasks. We address these gaps with BeyondMimic, a real-world framework to learn from human motions for versatile and naturalistic humanoid control via guided diffusion. Our framework provides a motion tracking pipeline capable of challenging skills such as jumping spins, sprinting, and cartwheels with state-of-the-art motion quality. Moving beyond simply mimicking existing motions, we further introduce a unified diffusion policy that enables zero-shot task-specific control at test time using simple cost functions. Deployed on hardware, BeyondMimic performs diverse tasks at test time, including waypoint navigation, joystick teleoperation, and obstacle avoidance, bridging sim-to-real motion tracking and flexible synthesis of human motion primitives for whole-body control. https://beyondmimic.github.io/.
@@ I Introduction
Recent advances in physics-based character animation [1, 2, 3, 4, 5] have already shown incredible progress to synthesize human motions into dynamic behaviors for whole-body control to solve downstream tasks. However, these successes remain limited to simulation, where agents enjoy idealized dynamics, unlimited actuation, and perfect observations, unlike real-world humanoids that face unmodeled dynamics, hardware limits, and imperfect state estimation. Achieving similar capabilities on hardware requires two missing capabilities: (1) a scalable, high-quality motion tracking framework that transforms kinematic references into robust, highly-dynamic motions while avoiding over-randomization, jitter, and degraded motion quality seen in prior work, and (2) an effective sim-to-real recipe to distill learned motion primitives into a single policy capable of composing skills at test time for flexible, goal-driven control without retraining.
@@
We introduce BeyondMimic, a real-world framework to address the above two challenges altogether. BeyondMimic first introduces a motion tracking pipeline that is able to perform highly dynamic motions, such as jumping spins, sprinting, and cartwheels, on real humanoid hardware with state-of-the-art motion quality. Beyond simply mimicking these human motions on humanoids, our framework further presents a unified guided-diffusion policy that synthesizes these motions or skills into novel motions for zero-shot, task-specific control at test time using only simple cost functions, solving downstream tasks flexibly without additional retraining.
@@
  • Guided Diffusion for Humanoids: Demonstrating loss-guided diffusion for real-world humanoid whole-body control, enabling diverse downstream tasks via simple loss functions at test time, a capability not previously achievable with existing methods..
@@ II-A Motion Tracking
Earlier motion-tracking approaches typically rely on a small set of similar reference motions to train single-task policies. These policies are co-trained with task-specific rewards alongside the DeepMimic objective, producing controllers capable of performing one task in one style. Examples include quadrupedal locomotion [11] and goalkeeping [12], and bipedal jumping and running [13]. To enable more dynamic skills, ASAP [14] proposes to use a real-to-sim pipeline that learns a delta action model from hardware experiments, improving simulation fidelity. However, this approach requires training motion-specific delta action models, which tend to overfit to the exact short motion. Similarly, KungfuBot [15] and HuB [16] achieve high-quality sim-to-real transfer with carefully hand-crafted domain randomization, but only for single short motions.
To move beyond single-motion policies, recent research explores scalable motion-tracking frameworks that can learn a diverse set of motions within a single policy. In physics-based animation, PHC [17] is a leading example of such a system, inspiring subsequent attempts in robotics to build general-purpose motion trackers. Early multi-motion trackers in robotics, such as OmniH2O [18], Exbody [19], and HumanPlus [20], demonstrate the feasibility of this approach but suffer significant motion quality degradation compared to their graphics counterparts. More recently, TWIST [21] enables good quality motion tracking but primarily for static motions, and similarly, CLONE [22] and UniTracker [23] focus mainly on low-dynamic walking motions with responsive upper-body tracking. On the other hand, GMT [24] handles some dynamic motions but sacrifices global trajectory tracking in favor of relative velocity tracking and gait regularization reward to improve robustness. Grandia et al. [25] present perhaps the highest-quality multi-motion tracking to date, but only on a small robot and without dynamic motions.
@@ II-B Diffusion in Robotics and Character Animation
To avoid the planning-control gap, another line of work focuses on learning a single, end-to-end policy that maps directly from state to an action distribution. This approach, Diffusion Policy [31], extends naturally to locomotion and character control. For example, DiffuseLoco [32], and BiRoDiff [33] learn unified policies that encode a wide range of skills, achieving smooth gait transitions and robust real-world deployment for quadruped robots. Meanwhile, PDP [34] extends the framework to physics-based animation, showing text prompt following and multi-modal behavior in push recovery. Using diffusion or flow matching policies in RL training also shows improved overall return and efficiency in motion tracking tasks [35]. While effective, such policies lack a mechanism for flexible test-time conditioning, such as loss-guided diffusion. This is because there is no straightforward way to compare a task goal defined in state space with an action sequence defined in joint space. As a result, conditioning on new tasks typically requires retraining.
@@ III Scaleable Motion Tracking
In this section, we detail the scalable pipeline for training motion-tracking policies. Using the same MDP and hyperparameters, the pipeline produces high-quality sim-to-real policies from minutes-long reference motions, demonstrating scalability across diverse motions. Note that since our tracking framework does not have a history, we omit timestep $t$ in the formulation for clarity. We use the subscript m to denote quantities from the reference motion. Unless otherwise stated, all amounts in this section are expressed in the world frame.
@@ III-B Observations
We use asymmetric actor–critics to boost training efficiency. In addition to the policy observation, the critic also receives per-body relative poses w.r.t. the anchor, $T^{-1}_{b_{\text{anchor}}}\,T_{b},\ \forall b\in\mathcal{B}$, enabling it to estimate body-wise tracking errors directly in Cartesian space.
@@ III-D Rewards
Task rewards are body tracking rewards. First, we compute error metrics for each target body $b\in\mathcal{B}_{\text{target}}$ based on the desired $(\hat{T}_{b},\hat{\mathcal{V}}_{b})$ and actual $(T_{b},\mathcal{V}_{b})$ poses and twists: $\mathbf{e}_{p,b}=\hat{\mathbf{p}}_{b}-\mathbf{p}_{b}$, $\mathbf{e}_{R,b}=\log(\hat{R}_{b}R_{b}^{\top})$, $\mathbf{e}_{v,b}=\hat{\mathbf{v}}_{b}-\mathbf{v}_{b}$, and $\mathbf{e}_{w,b}\approx\hat{\mathbf{w}}_{b}-\mathbf{w}_{b}$ assuming the orientaion error is small. Then, the mean squared errors are computed across all target bodies: $\bar{e}_{\chi}=\frac{1}{|\mathcal{B}_{\text{target}}|}\sum_{b\in\mathcal{B}_{\text{target}}}\|\mathbf{e}_{\chi,b}\|^{2},\,\chi\in\{p,R,v,w\}$. Each error metric is then normalized using a Gaussian-style exponential function: $r(\bar{e}_{\chi},\sigma)=\exp\left(-{\bar{e}_{\chi}}/{\sigma^{2}_{\chi}}\right)$, where $\sigma_{\chi}$ is a nominal error determined empirically.
The combined task reward is then defined as:
$r_{\text{task}}=\sum_{\chi\in\{p,R,v,w\}}r\!\bigl{(}\bar{e}_{\chi},\sigma_{\chi}\bigr{)}.$ (1)
For regularization, we include as few as three penalties critical for sim-to-real alignment. Among them, the joint limit penalty, $r_{\text{limit}}$, encourages joint positions to remain within the soft limits. The action rate penalty, $r_{\text{smooth}}$, encourages consecutive actions to be smooth, avoiding policies with excessive jitter. To penalize self-collision, we count the number of bodies where the self-contact force exceeds a predefined threshold as the total penalty, $r_{\text{contact}}$, over the bodies $b\notin\mathcal{B}_{\text{ee}}\subseteq\mathcal{B}$, where $\mathcal{B}_{\text{ee}}$ is the end-effector body set.
@@
$r=r_{\text{task}}-\lambda_{l}r_{\text{limit}}-\lambda_{s}r_{\text{smooth}}-\lambda_{c}r_{\text{contact}}$ (2)
@@ III-F Adaptive Sampling
Since failures are more likely caused by suboptimal actions taken shortly before termination, we apply a non-causal convolution with an exponentially decaying kernel $k(u)=\gamma^{u}$, where $\gamma$ is the decay rate, to assign greater weight to recent past failures. The final sampling probability from Bin s is then:
$p_{\text{s}}=\frac{\sum_{u=0}^{K-1}\alpha^{u}\,\bar{r}_{\text{s}+u}}{\sum_{j=1}^{\text{S}}\sum_{u=0}^{K-1}\alpha^{u}\,\bar{r}_{j+u}},$ (3)
where $\bar{r}_{\text{s}}$ is the smoothed failure rate of Bin s. We further mix the probability $p_{\text{s}}$ with a uniform distribution to preserve coverage of easier bins and mitigate catastrophic forgetting, using $p_{\text{s}}^{\prime}=\lambda\frac{1}{B}+(1-\lambda)p_{\text{s}}$, where $\lambda$ is the uniform sampling ratio. Starting bins are then drawn from $\mathrm{Multinomial}(p_{1}^{\prime},\dots,p_{\text{S}}^{\prime})$, prioritizing challenging regions.
@@ IV-A Training
The training itself follows a standard denoising diffusion process. First, in a forward pass, we incrementally add Gaussian noise to ground-truth trajectories. A denoising network, $x_{0,\theta}$, then learns to reverse this corruption. It takes a noised trajectory $\bm{\tau}_{t}^{\bm{k}}$ as input and is trained to predict the original, clean trajectory $\hat{\bm{\tau}}_{t}^{0}=x_{0,\theta}(\bm{\tau}_{t}^{\bm{k}},\bm{O}_{t},\bm{k})$. The network’s parameters are optimized by minimizing the Mean Squared Error (MSE) loss against the ground truth:
$\mathcal{L}=\text{MSE}(x_{0,\theta}(\bm{\tau}_{t}^{\bm{k}},\bm{O}_{t},\bm{k}),\bm{\tau}_{t})$ (4)
@@ IV-B Guidance
$\nabla_{\bm{\tau}}\log p(\bm{\tau}\mid\bm{\tau}^{*})=\nabla_{\bm{\tau}}\log p(\bm{\tau})+\nabla_{\bm{\tau}}\log p(\bm{\tau}^{*}\mid\bm{\tau}),$ (6)
where $\bm{\tau}^{*}$ is the optimal trajectory. In our application, this guidance term is a differentiable, task-specific cost function, $G_{c}(\bm{\tau})$. By establishing a relationship between the cost and the conditional likelihood, $p(\bm{\tau}^{*}\mid\bm{\tau})\propto\exp(-G_{\text{c}}(\bm{\tau}))$, the guidance term simplifies to the negative gradient of the cost: $\nabla_{\bm{\tau}}G_{\text{c}}(\bm{\tau})$.
@@ IV-C Downstream Tasks
$G_{\text{js}}(\bm{\tau})=\frac{1}{2}\sum_{t^{\prime}=t}^{t+H}\|V_{xy,t^{\prime}}(\bm{\tau}_{t^{\prime}})-g_{v}\|^{2},$ (7)
where $\bm{\tau}$ is the predicted trajectory at timestep $t$, $V_{xy,t}$ pulls out the planar root velocity at timestep $t$ and $g_{v}\in\mathbb{R}^{2}$ is the goal root velocity from the joystick controller.
@@
$G_{\text{wp}}(\bm{\tau})=\sum_{t^{\prime}=t}^{t+H}(1-e^{-2d})\|P_{x}(\bm{s}_{t^{\prime}})-g_{p}\|^{2}+e^{-2d}\|V_{x,t^{\prime}}(\bm{\tau}_{t^{\prime}})\|^{2},$ (8)
@@
$G_{\text{sdf}}(\bm{\tau})=\sum_{t^{\prime}=t}^{t+H}\sum_{b\in\mathcal{B}}B(\text{SDF}(\mathbf{P}_{b,t^{\prime}}(\bm{\tau}))-r_{i},\delta),$ (9)
@@ V Motion Tracking Experiments and Results
This section presents the experimental validation of the motion tracking framework through both simulation and real-world deployment on the Unitree G1 humanoid robot, showing high-quality, robust and highly dynamic motions. Note that we train all policies with the same set of hyperparameters throughout this section.
@@ V-C Real-world Performance
TABLE II: Ablation on Iterations to Convergence with Adaptive Sampling (AS)
@@ VI-D State Representation Ablation
    • Local States: Cartesian positions ($\mathbb{R}^{3\mathcal{B}}$), and linear velocities ($\mathbb{R}^{3\mathcal{B}}$) of all bodies, expressed in the character frame at each timestep.
@@
    • Local States: Joint angles ($\mathbb{R}^{J}$), where $J$ is the number of joints, and joint velocities ($\mathbb{R}^{J}$).
@@
Body-Pos State 100% 80%
Joint-Rot State 72% 0%
Table III depicts the performance of each representation for each of the two evaluated tasks. We find that the Body-Pos state representation significantly outperforms Joint-Pos across both tasks. While Joint-Rot is theoretically more Markovian, small errors in joint position estimates accumulate through the kinematic chain, making this representation more susceptible to compounding diffusion prediction errors compared to directly predicting Cartesian body poses. Furthermore, adding guidance for joystick control on the Joint-Rot state fails quickly and abruptly, even after careful tuning, likely because the base controller lacks robustness, and guidance pushes it further out of distribution.